you are the Director at XYZ hospital. Your unit cost for ensure is $3.00 The fixed cost is $5000.

November, you had an average of 50 patients/d who received 1 can ensure BID. Total cost was____

on December, you had an average of 47 patients/d who received 1 can ensure BID. Total cost was____
Answer #1

Total Cost = Fixed cost + variable cost

Variable cost is given to be $3 per unit.

(Here 1 unit is taken as 1 can ensure BID).

______________________________________________

November: Average of 50 patients each recieving 1 can ensure BID or 1 unit

Therefore for each patient, it costs $3. Hence for 50 patients, it costs: 50*3 = $150. This is the variable cost. The fixed cost is given to be $5000. Hence Total cost =

=Total Cost = Fixed cost + variable cost

=> Total cost = 5000 + 150 = $5150

Hence total cost was $5150 is the answer.

_____________________________________________________

December: Average number of patients = 47

Per patient cost for ensure = $3. Hence for 47 patients, the cost becomes: 47*3 = $141. This is the variable cost. Fixed cost is given to be $5000. Hence total cost =

Total cost = fixed cost + variable cost

=> Total cost = 5000 + 141 = $5141

Hence for the month of December, total cost was $5141 which is the answer.